OpenStudy (anonymous):

sin^2 2x + sin^2 4x = sin^2 6x help me!!!

OpenStudy (anonymous):

What's the question?

OpenStudy (anonymous):

Is that the question ^^

OpenStudy (anonymous):

\[\sin2x= 2sinxcosx\]\[\sin^2x = (sinx)^2\]

OpenStudy (shubhamsrg):

well i didnt solve it but as you should easily see,, x=0 is one solution ! infact x =n(pi) is a solution

OpenStudy (shubhamsrg):

even x = npi/2 is a soln..

OpenStudy (anonymous):

the solutions are sin(2.x)=0;cos(4.x)=-1;cos(4.x)=1/2

OpenStudy (anonymous):

\[x=k.\pi/2;x=(k.\pi/2) \pm \pi/12;x=(2k+1)\pi/4\]

OpenStudy (anonymous):

\[K \epsilon Z\]